Make an estimate of the electrical conductivity of Potassium at 0°c if its thermal conductivity has a value of 102.5 W/m/K.
1
Expert's answer
2019-09-30T10:17:42-0400

The Wiedemann–Franz law states


"\frac{\kappa}{\sigma}=LT"

where "\kappa" is the thermal conductivity, "\sigma" is the electrical conductivity, "T" is the temperature.

The Lorenz constant is equal

"L=2.44\times 10^{-8}\:\rm W\:\Omega\: K^{-2}"

Therefore, the electrical conductivity of Potassium at "\rm 0^{\circ}C"

"\sigma=\frac{\kappa}{LT}"

"=\frac{102.5\:\rm W/(m\cdot K)}{2.44\times 10^{-8}\:\rm W\:\Omega\: K^{-2}\times 273\:\rm K}=1.58\times 10^{7}\:\rm S/m"